Subject: Notice of RAITH off line status at some time next week to replace the X motor on the LIS Stage on the RAITH. day and time to be determined...
From: James Conway <jwc@snf.stanford.edu>
Date: Wed, 14 May 2008 18:27:26 -0700
Wed, 14 May 2008 18:27:26 -0700





Good Afternoon Raith Community:

The Raith system is up and Users are on the tool this evening.

Current Issues and system status:

1. *X Motor controller errors -- motor is going bad slowly.*

I was able to make changes to the X motor speed settings, reducing the 
speed of travel during movements of the Laser Interferometer Stage, in 
order to eliminate for a time the X motor control errors that we 
encountered yesterday evening.  This is just a short term fix as the 
motor is slowly failing.  A new motor is on order and it will be 
installed upon its arrival.  At some point next week likely Tuesday or 
Wednesday afternoon I will take the system away from scheduled users to 
perform this work.  Expect the system to be off line 8 - 10 hours before 
resuming the schedule.

2. *Neslab Chiller and cooling loop service scheduled for **this **Friday.*

The situation is deteriorating rapidly with reduced flow through the 
cooling water loop and likely a restriction through the upper turbo pump 
cooling jacket from corrosion or debris in the lines as in the past.
It is now less than 10 % of specified flow rate and I am starting to see 
some effects on the SEM platform such as drifting of the SEM image 
during high magnification imaging.  We also had a possible arcing event 
last evening as the Column Chamber Valve was actuated during a Load Lock 
procedure.  The EHT power supply and the FE-Gun tripped off and I was 
able to quickly recover the beam and bring up the system again for the 
next User.

So far we have not yet had a water flow or temperature alarm situation.  
Currently the E/O card temperature is 30.3 degrees C. and rising 
slowly.  I give it a 50:50 chance we will be able to continue operations 
through Friday at 10 AM when I take the system off-line for these 
repairs.  We have  a plan in place to have the system back up and ready 
for Users by about 4 PM Friday.

3.  *Universal Sample Holder and Standard Sample Holder measurements 
have been completed.*  No further changes to the Piezo post heights 
appear to be called for examining my initial data sets but I will be 
repeating these measurements and examining the total Delta WD 
measurements across a 90 mm transit across the 100 mm wafers.

Currently for a clean wafer mounted on the _Universal Sample Holder_ I 
was able to level the wafer successfully using the leveling routine. You 
can level to within +-2.71 um across all three points using the leveling 
routine. I was able to repeat this measurement three times on three 
different wafers with nearly identical results.

Currently for a clean 100 mm wafer mounted on the _Standard Sample 
Holder_, clipping with clips 2 and 4 inner positions , I was able to 
level the wafer successfully using the stage leveling routine on two 
runs to within +- 2.61 Áms after leveling.  I was able to perform this 
two times with the same wafer.  On the third attempt to level the same 
wafer we tested today I was unable to level within the range of the 
piezo post 100 Ám limit of travel.  The measurements indicated a 90 Ám 
total delta WD during the leveling routine.  Likely this was a particle 
on the wafer backside or a new scratch on the sample holder.  We plan to 
repeat these measurements in coming weeks to confirm and obtain better 
statistics.

The _ElectroStatic Chuck (ESC) _has not be tested yet but I have writes 
in the plan for early June which will test this fixture and we will make 
mechanical adjustments to this holder as necessary.

4. Please welcome Lev Markov...
Finally I would like to welcome the newest addition to RAITH USA Inc.,  
Field Service Engineer Lev Markov.  Lev Markov will be living, working 
and based in the SF Bay area to improve the technical and maintenance 
support for our RAITH 150 tool and will quickly become a very valuable 
resource for the Ebeam Lab.  He will be working with me from 10 - 4 PM 
Friday so if you are in the Cleanroom, please visit the Ebeam Lab and 
assist me in welcoming him here to the Bay Area.  Lev Markov brings with 
him a wealth of experience and many skills to the RAITH USA Inc. 
organization and will become our primary service contact.


*All Users are asked to be careful with the system and to report any 
unusual error messages during their operations to me by email.*

I will be out of the office tomorrow Thursday visiting UC Berkeley's 
Nanofabrication Labs.  SMS is preferred way to reach me during 9:30 AM - 
5 PM tomorrow.

Thank you for your support!

James Conway








["multipart/related" not shown]





Good Afternoon Raith Community:

The Raith system is up and Users are on the tool this evening.

Current Issues and system status:

1. X Motor controller errors -- motor is going bad slowly.

I was able to make changes to the X motor speed settings, reducing the speed of travel during movements of the Laser Interferometer Stage, in order to eliminate for a time the X motor control errors that we encountered yesterday evening.  This is just a short term fix as the motor is slowly failing.  A new motor is on order and it will be installed upon its arrival.  At some point next week likely Tuesday or Wednesday afternoon I will take the system away from scheduled users to perform this work.  Expect the system to be off line 8 - 10 hours before resuming the schedule.

2. Neslab Chiller and cooling loop service scheduled for
this Friday.

The situation is deteriorating rapidly with reduced flow through the cooling water loop and likely a restriction through the upper turbo pump cooling jacket from corrosion or debris in the lines as in the past.
It is now less than 10 % of specified flow rate and I am starting to see some effects on the SEM platform such as drifting of the SEM image during high magnification imaging.  We also had a possible arcing event last evening as the Column Chamber Valve was actuated during a Load Lock procedure.  The EHT power supply and the FE-Gun tripped off and I was able to quickly recover the beam and bring up the system again for the next User.

So far we have not yet had a water flow or temperature alarm situation.  Currently the E/O card temperature is 30.3 degrees C. and rising slowly.  I give it a 50:50 chance we will be able to continue operations through Friday at 10 AM when I take the system off-line for these repairs.  We have  a plan in place to have the system back up and ready for Users by about 4 PM Friday.

3.  Universal Sample Holder and Standard Sample Holder measurements have been completed.  No further changes to the Piezo post heights appear to be called for examining my initial data sets but I will be repeating these measurements and examining the total Delta WD measurements across a 90 mm transit across the 100 mm wafers.

Currently for a clean wafer mounted on the Universal Sample Holder
I was able to level the wafer successfully using the leveling routine. You can level to within +-2.71 um across all three points using the leveling routine. I was able to repeat this measurement three times on three different wafers with nearly identical results.

Currently for a clean 100 mm wafer mounted on the Standard Sample Holder, clipping with clips 2 and 4 inner positions , I was able to level the wafer successfully using the stage leveling routine on two runs to within +- 2.61 µms after leveling.  I was able to perform this two times with the same wafer.  On the third attempt to level the same wafer we tested today I was unable to level within the range of the piezo post 100 µm limit of travel.  The measurements indicated a 90 µm total delta WD during the leveling routine.  Likely this was a particle on the wafer backside or a new scratch on the sample holder.  We plan to repeat these measurements in coming weeks to confirm and obtain better statistics.

The ElectroStatic Chuck (ESC) has not be tested yet but I have writes in the plan for early June which will test this fixture and we will make mechanical adjustments to this holder as necessary.

4. Please welcome Lev Markov...
Finally I would like to welcome the newest addition to RAITH USA Inc.,  Field Service Engineer Lev Markov.  Lev Markov will be living, working and based in the SF Bay area to improve the technical and maintenance support for our RAITH 150 tool and will quickly become a very valuable resource for the Ebeam Lab.  He will be working with me from 10 - 4 PM Friday so if you are in the Cleanroom, please visit the Ebeam Lab and assist me in welcoming him here to the Bay Area.  Lev Markov brings with him a wealth of experience and many skills to the RAITH USA Inc. organization and will become our primary service contact.


All Users are asked to be careful with the system and to report any unusual error messages during their operations to me by email.

I will be out of the office tomorrow Thursday visiting UC Berkeley's Nanofabrication Labs.  SMS is preferred way to reach me during 9:30 AM - 5 PM tomorrow.

Thank you for your support!

James Conway







["image/gif" not shown]